OpenStudy (anonymous):

HELP FOR A MEDAL Justify if completing the square is a good method for solving when the Discriminant is negative. Use any of your three functions as an example and respond in complete sentences.

OpenStudy (anonymous):

f(x)=x2+6x+6 g(x)= x2-2 h(x)= f(x) = x2 + 1

OpenStudy (helder_edwin):

the quadratic formula for solving a quadratic equation comes from completing the square in a general quadratic equation. so both methods are pretty much the same.

OpenStudy (anonymous):

Im confused :/

OpenStudy (wolf1728):

I'd say using the quadratic formula is better than completing the square.

OpenStudy (anonymous):

but then why isnt the square a good method ?

OpenStudy (anonymous):

for solving when the Discriminant is negative.

OpenStudy (wolf1728):

I don't know why completing the square is better when the discriminant is negative.

OpenStudy (anonymous):

Meh :// okay thanks

OpenStudy (wolf1728):

Anyway, on the first half of this page, http://www.1728.org/quadr2.htm it talks about completing the square. Would you rather use this instead of the quadratic formula? I wouldn't.

OpenStudy (anonymous):

Nope nope, thanks so much c:

OpenStudy (wolf1728):

Glad I could help in some way :-)